OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = −3x + 12 and passes through the point (−1, 6). y = one thirdx + 7 y = −3x + 3 y = one thirdx + 3 y = −3x + 7

OpenStudy (anonymous):

y = −3x + 12 can you find the slope of this line??

OpenStudy (anonymous):

???

OpenStudy (anonymous):

y = mx + b

OpenStudy (anonymous):

yes here m is the slope now tell me the slope from the given equation

OpenStudy (anonymous):

-3?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

now when two lines are parallel their slopes are equal thus you got the slope for your line which is???

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

yes now you have slope = m = -3 and one point (-1,6) use point slope from to get the equation

OpenStudy (anonymous):

point slope form is y-y1 =m(x-x1)

OpenStudy (anonymous):

Im confused...

OpenStudy (anonymous):

ok look here

OpenStudy (anonymous):

you have slope = -3 right and one point mentioned in your question (-1,6) right???

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

\'''''

OpenStudy (anonymous):

using point slope form to get the equation y-y1 =m(x-x1) is a point slope form of line here y1 =6 x1 = -1 and m = -3 can you find the equation now???

OpenStudy (anonymous):

-1 = -18

OpenStudy (anonymous):

?

OpenStudy (anonymous):

no no

OpenStudy (anonymous):

we have y-y1 =m(x-x1) y-6 =-3(x-(-1)) y-6 = -3(x+1) do it further

OpenStudy (anonymous):

y - 6 = -3x + -3

OpenStudy (anonymous):

add 6 to both sides what you get???

OpenStudy (anonymous):

3

OpenStudy (anonymous):

y = -3x+3 is your answer

OpenStudy (anonymous):

Yay! Thank you!

OpenStudy (anonymous):

welcome